[Brahms-l] Shift report 20050314 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Mar 14 2005 - 23:57:18 EST
Shifter leader was RD
Shifters were 


1.  PHOBOS magnet problems during a fraction of the shift. Lost two stores.
2.  FS at 6 deg  1/4B collected ~500K trig 6 and will reach the 1.5M in the next shift.
3.  No new detector problems detected.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Mar 15 00:20:31 2005

This archive was generated by hypermail 2.1.8 : Tue Mar 15 2005 - 00:20:35 EST